# Benchmarking LLMs via Uncertainty Quantification

### v1

@@ 6.2 Evaluated Models
We select a diverse set of eight representative models (or model series) from the vast array of open-source LLMs available. These models encompass various architectures and training methodologies, allowing for a comprehensive benchmarking analysis. Specifically, the chosen models include the Llama-2 series (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Falcon series [^4] We omit Falcon-180B due to insufficient GPU resources. (Almazrouei et al., 2023), MPT-7B (Team, 2023b), Qwen series (Bai et al., 2023), Yi series (01.AI, 2023), DeepSeek series (DeepSeek, 2023), and InternLM-7B (Team, 2023a). For all models, we utilize their checkpoints from the HuggingFace platform [^5] https://huggingface.co/models.
@@ 6.3 Main Findings
In our primary experiments, we focus on LLMs with sizes ranging from 6B to 14B parameters. The outcomes are presented in Table 1. In addition to the three evaluation metrics, we also report the coverage rate, which is defined as the proportion of test instances for which the prediction set encompasses the true label. As previously mentioned, the reported results are the mean value derived from the two conformal score functions, LAC and APS. For a detailed analysis of the results pertaining to each function, please refer to Appendix C.1.
@@
In principle, an LLM having higher accuracy should exhibit lower uncertainty. However, the results regarding the SS metric reveal that in practice, higher accuracy does not necessarily correlate with lower uncertainty. Concretely, for each task, we observe that the ranking of LLMs based on accuracy differs from that based on uncertainty, suggesting that some LLMs possessing higher accuracy actually display higher uncertainty. Notably, two LLMs with a significant difference in accuracy may even display inverse uncertainty. For example, on the DRS task, InternLM-7B demonstrates higher accuracy than MPT-7B in accuracy by 19.34 absolute points, yet it shows higher uncertainty. This pattern is also observed on the QA task with Qwen-7B and Llama-2-7B (9.61), the CI task with Qwen-14B and Yi-6B (14.50), and the DS task with InternLM-7B and Falcon-7B (9.69). In each case, the LLM with much higher accuracy exhibits greater uncertainty compared to its counterpart. These observations underscore the importance of considering uncertainty in addition to accuracy when evaluating LLMs.
Our new metric, UAcc, takes into account both accuracy and uncertainty. It can penalize LLMs with high uncertainty and reward those with low uncertainty, as exemplified by Qwen-14B on the QA task and the RC task, respectively. This unique characteristic of UAcc allows it to enlarge or shrink the relative improvement of one LLM compared to another. For example, on the QA task, Qwen-14B demonstrates an 11.60% higher accuracy than Yi-6B, but when considering UAcc, the improvement increases to 28%. Conversely, on the CI task, InternLM-7B has an 8.32% higher accuracy than DeepSeek-7B, but the improvement in UAcc is only 0.96%. Moreover, the UAcc metric can even alter the relative ranking of two LLMs. For example, on the DS task, Llama-2-13B achieves lower accuracy but higher UAcc than Mistral-7B. These results highlight the unique contribution of UAcc in providing a more nuanced evaluation of LLMs.
@@ 6.4 Effects of Model Scale
LLMs with larger sizes are usually pretrained on more data and tend to exhibit superior capabilities across various tasks. In this study, we aim to investigate how an LLM’s performance changes when scaling its model size. We present the results of the Qwen model series in Figure 3. It is evident that, with the exception of Qwen-72B and Qwen-14B on the CI task, an increase in model size consistently leads to improved task performance. Concerning uncertainty, there is a general trend of decreasing uncertainty when scaling the model size from 1.8B to 14B. However, Qwen-7B displays higher uncertainty compared to Qwen-1.8B on the QA task.
When further scaling the model size from 14B to 72B, the enhancements in uncertainty become less pronounced, and more variations are observed. Notably, on both the RC and DRS tasks, Qwen-72B demonstrates higher uncertainty than Qwen-14B, although Qwen-72B achieves higher accuracy. Consequently, Qwen-72B underperforms Qwen-14B in terms of UAcc on the two tasks. We provide the results of Llama-2 series, Yi series, DeepSeek series, and Falcon series in Appendix C.2, which reveal similar findings.
@@ 6.5 Effects of Instruction Finetuning
Figure 4 shows that Chat-V1 consistently results in lower accuracy and higher uncertainty across all model sizes. Conversely, Chat-V2 enhances accuracy for Llama-2-7B and Llama-2-13B. However, for Llama-2-70B, Chat-V2 also leads to a decline in accuracy. Regarding uncertainty, Chat-V2 consistently results in increased uncertainty compared to the base model, although the extent of degradation is less severe than Chat-V1. With respect to UAcc, both Chat-V1 and Chat-V2 exhibit inferior performance to the base model. These findings suggest that instruction-finetuning tends to impair model performance, particularly in terms of uncertainty.
@@ 6.6 Effects of Amount of Calibration Data
As described in Section 3, conformal prediction requires a calibration set to calculate the threshold $\hat{q}$. In our prior analyses, we have allocated 50% of the data as the calibration set and the remaining 50% as the test set. Here, we explore the impact of varying the proportion of calibration data, ranging from 10% to 50%, on uncertainty quantification. Note that the same 50% of data is consistently used as the test set. The average results across five tasks of the Llama-2 series are shown in Figure 5. It is observed that there are no significant variations in coverage rate, uncertainty (SS), and UAcc for all versions of the Llama-2 series when varying the amount of calibration data. This observation confirms the efficacy of applying conformal prediction for uncertainty quantification in our analysis.
LLMs Acc (%) $\uparrow$ SS $\downarrow$ UAcc (%) $\uparrow$
Yi-34B 81.46 1.93 121.09
$*{}^{*}$ 80.10 1.88 119.99
Llama-2-70B 72.24 2.16 94.62
$*{}^{*}$ 63.55 2.72 63.67
Table 2: Average results across five tasks of Yi-34B and Llama-2-70B. *** indicates the results obtained without using demonstrations in the prompt input.
@@ 6.7 Ablation Study
Inspired by in-context learning (Xie et al., 2021), our previous analyses have incorporated demonstrations for each task. Here, we aim to compare the performance of LLMs when demonstrations are included or excluded from the prompt input. The results of Yi-34B and Llama-2-70B are reported in Table 2. We observe that the presence of demonstrations leads to improved performance for both models, as evidenced by increased accuracy and UAcc. Having demonstrations also leads to lower uncertainty for Llama-2-70B but higher uncertainty for Yi-34B. Overall, these results substantiate the effectiveness of incorporating demonstrations into the prompt to stimulate the capabilities of LLMs.
@@ 7 Conclusion
In this work, we have provided an extensive examination of the performance of LLMs by focusing on prediction uncertainty. To achieve this, we have employed conformal prediction for uncertainty quantification. Our comprehensive investigation, which involves eight LLMs (or LLM series) and spans five tasks, demonstrates that relying solely on accuracy for benchmarking LLMs is insufficient. Instead, it is imperative to take uncertainty into account when assessing their overall performance.
@@ Appendix A Pseudo Code
2:Evaluation results in terms of $Acc$, $SS$, $UAcc$;
@@
27:$UAcc\leftarrow\frac{Acc}{SS}\sqrt{|\mathcal{Y}|}$, where $\mathcal{Y}$ denotes the option set;
28:return $Acc$, $SS$, $UAcc$.
@@
    Calculate the evaluation metrics, namely Acc, SS, and UAcc.
@@ Appendix B Dataset Statistics
LLMs Coverage Rate (%) Acc (%) $\uparrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 90.43 91.57 91.40 90.31 89.63 90.67 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1)
Yi-6B 89.96 90.04 89.83 90.96 90.40 90.24 57.57(3) 85.99(2) 76.50(2) 58.72(3) 66.06(1) 68.97(2)
Mistral-7B 89.64 89.48 89.99 91.01 90.86 90.20 60.44(2) 81.94(4) 62.93(4) 53.21(4) 62.16(2) 64.14(3)
Llama-2-13B 90.06 89.91 90.32 90.60 90.46 90.27 52.52(5) 77.23(5) 59.66(5) 52.65(5) 60.05(3) 60.42(4)
Qwen-7B 90.54 89.80 89.88 90.50 90.00 90.14 55.21(4) 83.89(3) 63.70(3) 64.04(2) 32.53(8) 59.87(5)
InternLM-7B 89.38 90.33 89.96 90.61 90.62 90.18 48.37(6) 73.86(6) 46.21(6) 43.72(6) 34.38(7) 49.31(6)
Llama-2-7B 90.42 89.61 91.00 89.79 89.85 90.13 45.60(8) 65.79(7) 43.05(7) 32.61(8) 45.60(5) 46.53(7)
DeepSeek-7B 89.94 88.48 89.79 91.09 90.16 89.89 45.65(7) 65.39(8) 42.66(8) 33.50(7) 42.15(6) 45.87(8)
MPT-7B 90.24 90.77 90.19 90.89 89.48 90.31 29.49(9) 31.69(9) 25.50(9) 24.38(10) 24.86(9) 27.18(9)
Falcon-7B 90.06 89.94 89.77 90.72 90.60 90.22 23.75(10) 24.98(10) 24.91(10) 25.86(9) 24.69(10) 24.84(10)
LLMs SS $\downarrow$ UAcc (%) $\uparrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 2.39(2) 1.00(1) 1.01(1) 1.54(1) 2.26(4) 1.64(1) 66.53(1) 223.90(1) 220.76(1) 117.86(1) 53.49(4) 136.51(1)
Yi-6B 2.92(5) 1.12(2) 1.53(2) 2.74(5) 1.60(1) 1.98(2) 49.68(3) 187.72(2) 122.33(2) 52.99(3) 101.34(1) 102.81(2)
Mistral-7B 2.32(1) 1.26(4) 2.36(4) 2.62(4) 2.14(3) 2.14(3) 63.78(2) 159.84(4) 65.39(4) 49.75(5) 71.22(2) 82.00(3)
Llama-2-13B 2.73(3) 1.58(5) 2.58(5) 2.53(3) 2.12(2) 2.31(5) 47.18(5) 119.60(5) 56.59(5) 51.07(4) 69.66(3) 68.82(5)
Qwen-7B 2.82(4) 1.21(3) 2.02(3) 2.08(2) 2.93(5) 2.21(4) 48.34(4) 169.90(3) 77.19(3) 75.66(2) 27.20(7) 79.66(4)
InternLM-7B 3.23(8) 1.71(6) 3.17(7) 3.54(8) 4.43(10) 3.22(8) 36.69(6) 106.07(6) 35.74(6) 30.42(6) 19.03(8) 45.59(6)
Llama-2-7B 3.05(6) 2.20(7) 3.23(8) 3.25(6) 3.45(7) 3.03(6) 36.65(7) 73.36(7) 32.68(8) 24.60(7) 32.38(6) 39.93(7)
DeepSeek-7B 3.19(7) 2.50(8) 3.01(6) 3.38(7) 3.09(6) 3.03(6) 35.16(8) 64.09(8) 34.89(7) 24.26(8) 33.38(5) 38.36(8)
MPT-7B 3.54(9) 3.44(9) 3.60(9) 3.62(9) 3.63(8) 3.57(9) 20.39(9) 22.60(9) 17.37(9) 16.50(10) 16.77(9) 18.73(9)
Falcon-7B 3.92(10) 3.59(10) 3.64(10) 3.66(10) 3.93(9) 3.75(10) 14.85(10) 17.03(10) 16.76(10) 17.30(9) 15.38(10) 16.26(10)
Table 3: Results of LLMs with sizes ranging from 6B to 14B. These results are obtained when LAC is adopted as the conformal score function. The "Avg." column denotes the average performance across tasks. The small number in parentheses indicates the rank of the model on each task. Note that UAcc takes values in the range $[0,\sqrt{|\mathcal{Y}|}]$, thus can be larger than 1.
LLMs Coverage Rate (%) Acc (%) $\uparrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 94.74 98.83 99.19 93.52 89.79 95.21 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1)
Yi-6B 92.64 98.09 94.86 91.77 92.37 93.95 57.57(3) 85.99(2) 76.50(2) 58.72(3) 66.06(1) 68.97(2)
Mistral-7B 96.37 96.33 89.88 91.04 93.24 93.37 60.44(2) 81.94(4) 62.93(4) 53.21(4) 62.16(2) 64.14(3)
Llama-2-13B 95.12 96.39 90.68 91.25 91.18 92.92 52.52(5) 77.23(5) 59.66(5) 52.65(5) 60.05(3) 60.42(4)
Qwen-7B 95.04 98.25 93.18 94.35 89.12 93.99 55.21(4) 83.89(3) 63.70(3) 64.04(2) 32.53(8) 59.87(5)
InternLM-7B 91.98 96.23 90.24 90.18 90.06 91.74 48.37(6) 73.86(6) 46.21(6) 43.72(6) 34.38(7) 49.31(6)
Llama-2-7B 92.33 91.76 90.95 89.40 90.24 90.94 45.60(8) 65.79(7) 43.05(7) 32.61(8) 45.60(5) 46.53(7)
DeepSeek-7B 92.42 91.43 90.53 90.70 90.26 91.07 45.65(7) 65.39(8) 42.66(8) 33.50(7) 42.15(6) 45.87(8)
MPT-7B 89.35 90.32 90.06 90.72 89.94 90.08 29.49(9) 31.69(9) 25.50(9) 24.38(10) 24.86(9) 27.18(9)
Falcon-7B 90.01 89.96 89.86 90.20 90.82 90.17 23.75(10) 24.98(10) 24.91(10) 25.86(9) 24.69(10) 24.84(10)
LLMs SS $\downarrow$ UAcc (%) $\uparrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 3.21(1) 2.47(2) 3.03(5) 2.33(1) 2.47(3) 2.70(1) 49.12(1) 91.15(1) 73.50(2) 77.53(1) 48.95(4) 68.05(1)
Yi-6B 3.48(5) 2.72(5) 2.22(1) 2.95(4) 2.33(1) 2.74(3) 40.68(3) 77.50(3) 84.49(1) 48.95(4) 69.64(1) 64.25(2)
Mistral-7B 3.27(2) 2.25(1) 2.59(3) 2.80(3) 2.67(4) 2.71(2) 45.42(2) 89.57(2) 59.50(4) 46.62(5) 57.28(3) 59.68(3)
Llama-2-13B 3.40(4) 2.90(6) 2.86(4) 2.58(2) 2.36(2) 2.82(4) 37.88(4) 65.32(6) 51.06(5) 49.98(3) 62.38(2) 53.32(4)
Qwen-7B 3.70(8) 3.10(8) 2.53(2) 2.95(4) 2.91(5) 3.04(5) 36.56(5) 66.31(5) 61.74(3) 53.18(2) 27.36(7) 49.03(5)
InternLM-7B 3.74(9) 2.68(4) 3.39(8) 3.71(10) 4.51(10) 3.61(9) 31.65(8) 67.60(4) 33.38(7) 29.03(6) 18.71(8) 36.07(7)
Llama-2-7B 3.35(3) 2.58(3) 3.32(7) 3.27(6) 3.15(7) 3.14(6) 33.30(6) 62.48(7) 31.83(8) 24.40(7) 35.43(5) 37.49(6)
DeepSeek-7B 3.48(5) 3.03(7) 3.12(6) 3.42(7) 3.07(6) 3.23(7) 32.10(7) 52.91(8) 33.56(6) 23.97(8) 33.67(6) 35.24(8)
MPT-7B 3.53(7) 3.49(9) 3.60(9) 3.55(8) 3.69(8) 3.57(8) 20.49(9) 22.26(9) 17.35(9) 16.82(10) 16.49(9) 18.68(9)
Falcon-7B 3.89(10) 3.60(10) 3.69(10) 3.62(9) 3.91(9) 3.74(10) 14.96(10) 16.98(10) 16.56(10) 17.52(9) 15.47(10) 16.30(10)
Table 4: Results of LLMs with sizes ranging from 6B to 14B. These results are obtained when APS is adopted as the conformal score function. The "Avg." column denotes the average performance across tasks. The small number in parentheses indicates the rank of the model on each task. Note that UAcc takes values in the range $[0,\sqrt{|\mathcal{Y}|}]$, thus can be larger than 1.

### v2

@@ 6.2 Evaluated Models
We select a diverse set of nine representative models (or model series) from the vast array of open-source LLMs available. These models encompass various architectures and training methodologies, thereby allowing for a comprehensive benchmarking analysis. Specifically, the chosen models include the Llama-2 series (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Falcon series [^5] We omit Falcon-180B due to insufficient GPU resources. (Almazrouei et al., 2023), MPT-7B (Team, 2023b), Gemma-7B (Team et al., 2024), Qwen series (Bai et al., 2023), Yi series (01.AI, 2023), DeepSeek series (DeepSeek, 2023), and InternLM-7B (Team, 2023a). For all models, we utilize their checkpoints from the HuggingFace platform [^6] https://huggingface.co/models.
@@ 6.3 Main Findings
In our primary experiments, we focus on LLMs with sizes ranging from 6B to 14B parameters. The outcomes of CR, Acc and SS are presented in Tables 1 and 2, respectively. As previously mentioned, the reported results are the mean value derived from the two conformal score functions, LAC and APS. For a detailed analysis of the results pertaining to each score function, please refer to Appendix C.1.
@@
In principle, an LLM having higher accuracy should demonstrate lower uncertainty. However, as shown in Table 2, the results regarding the SS metric reveal that in practice, higher accuracy does not necessarily correlate with lower uncertainty. Concretely, for each task, we observe that the ranking of LLMs based on accuracy differs from that based on uncertainty, suggesting that some LLMs possessing higher accuracy actually display higher uncertainty. Notably, two LLMs with a significant difference in accuracy may even display inverse uncertainty. For example, on the DRS task, InternLM-7B demonstrates higher accuracy than MPT-7B in accuracy by 19.34 absolute points, yet it shows higher uncertainty. This pattern is also observed on the QA task with Qwen-7B and Llama-2-7B (9.61), the CI task with Qwen-14B and Yi-6B (14.50), and the DS task with InternLM-7B and Falcon-7B (9.69). In each case, the LLM with much higher accuracy exhibits greater uncertainty compared to its counterpart. These observations underscore the importance of considering uncertainty in addition to accuracy when evaluating LLMs.
@@ 6.4 Effects of Model Scale
LLMs with larger sizes are usually pretrained on more data and tend to exhibit superior capabilities across various tasks. In this study, we aim to investigate how an LLM’s performance changes when scaling its model size. We present the results of the Qwen model series in Figure 3. It is evident that in the majority of cases, the coverage guarantee requirement has been satisfied. In terms of Acc, with the exception of Qwen-72B and Qwen-14B on the CI task, an increase in model size consistently leads to improved performance. Concerning uncertainty (i.e. SS), there is a general trend of decreasing uncertainty when scaling the model size from 1.8B to 14B. However, Qwen-7B displays higher uncertainty compared to Qwen-1.8B on the QA task. When further scaling the model size from 14B to 72B, the enhancements in uncertainty become less pronounced, and more variations are observed. Notably, on both the RC and DRS tasks, Qwen-72B demonstrates higher uncertainty than Qwen-14B, although Qwen-72B achieves higher accuracy.
We provide the results of the Llama-2 series, Yi series, DeepSeek series, and Falcon series in Appendix C.2, which reveal similar findings.
@@ 6.5 Effects of Instruction Finetuning
Figure 4 shows that Chat-V1 consistently results in lower accuracy and higher uncertainty across all model sizes. Conversely, Chat-V2 enhances accuracy for Llama-2-7B and Llama-2-13B. However, for Llama-2-70B, Chat-V2 also leads to a decline in accuracy. Regarding uncertainty, Chat-V2 consistently results in increased uncertainty compared to the base model, although the extent of degradation is less severe than Chat-V1. These findings suggest that instruction-finetuning tends to impair model performance, particularly in terms of uncertainty.
@@
LLMs CR (%) Acc (%) $\uparrow$ SS $\downarrow$
Mixtral-8x7B 92.59 73.74 2.03
Mistral-7B 91.78 64.14 2.43
Llama-2-13B 91.60 60.42 2.56
Qwen-14B 92.94 74.00 2.17
Table 3: Average results across five tasks of Mixtral-8x7B, Mistral-7B, Llama-2-13B, and Qwen-14B.
6.6 Effects of Mixture of Experts
A Mixture of Experts (MoE) is a technique that combines multiple specialized models with a gating mechanism to enhance performance by leveraging diverse expertise and adaptability in handling complex data relationships. In recent months, the adoption of the MoE technique to augment the performance of LLMs has been steadily increasing. Considering this, we study how MoE impacts the uncertainty of LLMs, specifically comparing the Mixtral-8x7B MoE model [^9] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 with Mistral-7B. The results are reported in Table 3. While Mixtral-8x7B has 46.7B total parameters, it only uses 12.9B parameters per token. It, therefore, processes input and generates output at the same speed and for the same cost as a 12.9B model [^10] https://mistral.ai/news/mixtral-of-experts/. To provide a comprehensive comparison, we also include the results of Llama-2-13B and Qwen-14B. As can be observed, Mixtral-8x7B consistently outperforms Mistral-7B across both accuracy and uncertainty. Remarkably, it achieves accuracy levels comparable to Qwen-14B, while demonstrating the lowest average uncertainty among the four LLMs. These observations show that MoE is indeed an effective method to enhance the accuracy and reduce the uncertainty of LLMs.
@@ 6.7 Effects of Amount of Calibration Data
As described in Section 3, conformal prediction requires a calibration set to calculate the threshold $\hat{q}$. In our prior analyses, we have allocated 50% of the data as the calibration set and the remaining 50% as the test set. Here, we explore the impact of varying the proportion of calibration data, ranging from 10% to 50%, on uncertainty quantification. Note that the same 50% of data is consistently used as the test set. The average results across five tasks of the Llama-2 series are shown in Figure 5. It is observed that there are no significant variations in coverage rate (CR) and uncertainty (SS) for all versions of the Llama-2 series when varying the amount of calibration data. This observation confirms the efficacy of applying conformal prediction for uncertainty quantification in our analysis.
LLMs CR (%) Acc (%) $\uparrow$ SS $\downarrow$
Yi-34B 94.22 81.46 1.93
Yi-34B 93.31 80.10 1.88
Llama-2-70B 93.11 72.24 2.16
Llama-2-70B 91.55 63.55 2.72
Table 4: Average results across five tasks of Yi-34B and Llama-2-70B. * indicates the results obtained without using demonstrations in the prompt input.
@@ 6.8 Ablation Study
Inspired by in-context learning (Xie et al., 2021), our previous analyses have incorporated demonstrations for each task. Here, we aim to compare the performance of LLMs when demonstrations are included or excluded from the prompt input. The results of Yi-34B and Llama-2-70B are reported in Table 4. We observe that the presence of demonstrations leads to improved performance for both models, as evidenced by increased accuracy. Having demonstrations also leads to lower uncertainty for Llama-2-70B but higher uncertainty for Yi-34B. Overall, these results substantiate the effectiveness of incorporating demonstrations into the prompt to stimulate the capabilities of LLMs.
6.9 Comparison to Other Uncertainty Quantification Methods
Tasks CR (%) SS
CP PPL CP PPL
QA 90.68 83.44 3.49 2.89
RC 93.28 95.48 2.19 2.39
CI 90.10 96.25 3.28 3.97
DRS 90.40 86.80 3.63 3.42
DS 90.34 87.13 4.47 4.33
Avg. 90.96 89.82 3.41 3.40
Table 5: Comparison between conformal prediction (CP) and perplexity (PPL) using InternLM-7B.
Given the predicted probabilities, another widely used measure of uncertainty is entropy (Rényi, 1961). There have also been some entropy-based uncertainty quantification methods for language models (Fadeeva et al., 2023). Here, we compare conformal prediction to entropy. Since conformal prediction uses the prediction set size (SS) to measure uncertainty, a direct comparison with entropy is not straightforward. To address this issue, we convert entropy to perplexity (Jelinek et al., 1977), which is defined as $PPL=2^{H}$, where $H$ denotes the entropy. Perplexity takes values in the range of $[1,|\mathcal{Y}|]$, allowing it to be interpreted as prediction set size. For instance, when the predicted probability for each class (option) is $\frac{1}{|\mathcal{Y}|}$, $PPL=|\mathcal{Y}|$.
The results regarding InternLM-7B are presented in Table 5, from which, we observe that the coverage rate of perplexity varies significantly across different tasks. On the QA task, the coverage rate is only $83.44\%$. In contrast, the coverage rate of conformal prediction consistently exceeds $90\%$. These results demonstrate the advantages of using conformal prediction for uncertainty quantification. [^11] It is noted that except for coverage rate (CR) and prediction set size (SS), some other metrics such as AUROC (Kuhn et al., 2023) and PRR (Malinin and Gales, 2020) have been proposed to evaluate the performance of uncertainty quantification methods. However, it has been pointed out that these metrics have many limitations (Huang et al., 2024). Instead, both CR and SS are intuitive and highly interpretable, making them reasonable choices.
We provide more analyses in Appendix C.5.
@@ 7 Conclusion
In this work, we have provided an extensive examination of the performance of LLMs by focusing on prediction uncertainty. To achieve this, we have employed conformal prediction for uncertainty quantification. Our comprehensive investigation, which involves nine LLMs (or LLM series) and spans five tasks, demonstrates that relying solely on accuracy for benchmarking LLMs is insufficient. Instead, it is imperative to take uncertainty into account when assessing their overall performance.
@@ Appendix A Pseudo Code
2:Evaluation results in terms of $Acc$, $SS$, $CR$;
@@
27:$CR\leftarrow$ $\mathcal{B}$, $\mathcal{D}_{test}$);
28:return $Acc$, $SS$, $CR$.
@@
    Calculate the evaluation metrics, namely Acc, SS, and CR.
@@
LLMs LAC APS
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 90.43 91.57 91.40 90.31 89.63 90.67 94.74 98.83 99.19 93.52 89.79 95.21
Yi-6B 89.96 90.04 89.83 90.96 90.40 90.24 92.64 98.09 94.86 91.77 92.37 93.95
Gemma-7B 90.38 90.48 90.57 90.30 90.81 90.51 96.76 97.85 93.70 92.87 90.60 94.35
Mistral-7B 89.64 89.48 89.99 91.01 90.86 90.20 96.37 96.33 89.88 91.04 93.24 93.37
Llama-2-13B 90.06 89.91 90.32 90.60 90.46 90.27 95.12 96.39 90.68 91.25 91.18 92.92
Qwen-7B 90.54 89.80 89.88 90.50 90.00 90.14 95.04 98.25 93.18 94.35 89.12 93.99
InternLM-7B 89.38 90.33 89.96 90.61 90.62 90.18 91.98 96.23 90.24 90.18 90.06 91.74
Llama-2-7B 90.42 89.61 91.00 89.79 89.85 90.13 92.33 91.76 90.95 89.40 90.24 90.94
DeepSeek-7B 89.94 88.48 89.79 91.09 90.16 89.89 92.42 91.43 90.53 90.70 90.26 91.07
MPT-7B 90.24 90.77 90.19 90.89 89.48 90.31 89.35 90.32 90.06 90.72 89.94 90.08
Falcon-7B 90.06 89.94 89.77 90.72 90.60 90.22 90.01 89.96 89.86 90.20 90.82 90.17
Table 6: The CR (%) results of LLMs with sizes ranging from 6B to 14B. The "Avg." column denotes the average performance across tasks. The small number in parentheses indicates the rank of the model on each task.
LLMs Acc (%) $\uparrow$ SS (%) $\downarrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1) 2.39(3) 1.00(1) 1.01(1) 1.54(1) 2.26(4) 1.64(1)
Yi-6B 57.57(4) 85.99(2) 76.50(2) 58.72(4) 66.06(1) 68.97(2) 2.92(6) 1.12(2) 1.53(2) 2.74(6) 1.60(1) 1.98(2)
Gemma-7B 62.24(2) 85.29(3) 73.58(3) 66.79(2) 40.80(7) 65.74(3) 2.23(1) 1.18(3) 1.79(3) 1.97(2) 3.17(7) 2.07(3)
Mistral-7B 60.44(3) 81.94(5) 62.93(5) 53.21(5) 62.16(2) 64.14(4) 2.32(2) 1.26(5) 2.36(5) 2.62(5) 2.14(3) 2.14(4)
Llama-2-13B 52.52(6) 77.23(6) 59.66(6) 52.65(6) 60.05(3) 60.42(5) 2.73(4) 1.58(6) 2.58(6) 2.53(4) 2.12(2) 2.31(6)
Qwen-7B 55.21(5) 83.89(4) 63.70(4) 64.04(3) 32.53(9) 59.87(6) 2.82(5) 1.21(4) 2.02(4) 2.08(3) 2.93(5) 2.21(5)
InternLM-7B 48.37(7) 73.86(7) 46.21(7) 43.72(7) 34.38(8) 49.31(7) 3.23(9) 1.71(7) 3.17(8) 3.54(9) 4.43(11) 3.22(9)
Llama-2-7B 45.60(9) 65.79(8) 43.05(8) 32.61(9) 45.60(5) 46.53(8) 3.05(7) 2.20(8) 3.23(9) 3.25(7) 3.45(8) 3.03(7)
DeepSeek-7B 45.65(8) 65.39(9) 42.66(9) 33.50(8) 42.15(6) 45.87(9) 3.19(8) 2.50(9) 3.01(7) 3.38(8) 3.09(6) 3.03(7)
MPT-7B 29.49(10) 31.69(10) 25.50(10) 24.38(11) 24.86(10) 27.18(10) 3.54(10) 3.44(10) 3.60(10) 3.62(10) 3.63(9) 3.57(10)
Falcon-7B 23.75(11) 24.98(11) 24.91(11) 25.86(10) 24.69(11) 24.84(11) 3.92(11) 3.59(11) 3.64(11) 3.66(11) 3.93(10) 3.75(11)
Table 7: The Acc and SS results of LLMs with sizes ranging from 6B to 14B. These results are obtained when LAC is adopted as the conformal score function. The "Avg." column denotes the average performance across tasks. The small number in parentheses indicates the rank of the model on each task.
LLMs Acc (%) $\uparrow$ SS $\downarrow$
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Qwen-14B 64.25(1) 91.52(1) 91.00(1) 73.90(1) 49.33(4) 74.00(1) 3.21(1) 2.47(2) 3.03(6) 2.33(2) 2.47(3) 2.70(2)
Yi-6B 57.57(4) 85.99(2) 76.50(2) 58.72(4) 66.06(1) 68.97(2) 3.48(6) 2.72(6) 2.22(1) 2.95(5) 2.33(1) 2.74(4)
Gemma-7B 62.24(2) 85.29(3) 73.58(3) 66.79(2) 40.80(7) 65.74(3) 3.21(1) 2.57(3) 2.29(2) 2.31(1) 3.05(6) 2.68(1)
Mistral-7B 60.44(3) 81.94(5) 62.93(5) 53.21(5) 62.16(2) 64.14(4) 3.27(3) 2.25(1) 2.59(4) 2.80(4) 2.67(4) 2.71(3)
Llama-2-13B 52.52(6) 77.23(6) 59.66(6) 52.65(6) 60.05(3) 60.42(5) 3.40(5) 2.90(7) 2.86(5) 2.58(3) 2.36(2) 2.82(5)
Qwen-7B 55.21(5) 83.89(4) 63.70(4) 64.04(3) 32.53(9) 59.87(6) 3.70(9) 3.10(9) 2.53(3) 2.95(5) 2.91(5) 3.04(6)
InternLM-7B 48.37(7) 73.86(7) 46.21(7) 43.72(7) 34.38(8) 49.31(7) 3.74(10) 2.68(5) 3.39(9) 3.71(11) 4.51(11) 3.61(10)
Llama-2-7B 45.60(9) 65.79(8) 43.05(8) 32.61(9) 45.60(5) 46.53(8) 3.35(4) 2.58(4) 3.32(8) 3.27(7) 3.15(8) 3.14(7)
DeepSeek-7B 45.65(8) 65.39(9) 42.66(9) 33.50(8) 42.15(6) 45.87(9) 3.48(6) 3.03(8) 3.12(7) 3.42(8) 3.07(7) 3.23(8)
MPT-7B 29.49(10) 31.69(10) 25.50(10) 24.38(11) 24.86(10) 27.18(10) 3.53(8) 3.49(10) 3.60(10) 3.55(9) 3.69(9) 3.57(9)
Falcon-7B 23.75(11) 24.98(11) 24.91(11) 25.86(10) 24.69(11) 24.84(11) 3.89(11) 3.60(11) 3.69(11) 3.62(10) 3.91(10) 3.74(11)
Table 8: The Acc and SS results of LLMs with sizes ranging from 6B to 14B. These results are obtained when APS is adopted as the conformal score function. The "Avg." column denotes the average performance across tasks. The small number in parentheses indicates the rank of the model on each task.